Poisson distribution models rare events occurring at a constant average rate
λ = 3 represents the average number of accidents per day
The complement rule: P(N > 5) = 1 - P(N ≤ 5)
The answer 0.084 corresponds to option D
Get started today
Ultimate access to all questions.
If the rate at which accidents occur in a U.S. city is, on average, three every day, calculate the probability that more than 5 accidents occur on a single day.
Other
Community
NNikitesh
Last updated: February 2, 2026 at 10:22
0
A
0.15
33.3%
B
0.06
33.3%
C
0.05
D
0.084
33.3%
Powered ByGPT 5.6 Luna powered
Comments (0)
No comments yet.
If the rate at which accidents occur in a U.S. city is, on average, three every day, calculate the probability that more than 5 accidents occur on a single day. | Financial Risk Manager Part 1 Quiz - LeetQuiz